# Bird-Interact: Re-imagining Text-to-SQL Evaluation for Large Language Models via Lens of Dynamic Interactions

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance on single-turn text-to-SQL tasks, but real-world database applications predominantly require multi-turn interactions to handle ambiguous queries, execution errors, and evolving user requirements. Existing multi-turn benchmarks fall short of capturing this complexity, either by treating conversation histories as static context or by limiting evaluation to narrow, read-only (SELECT-ONLY) operations, thereby may potentially failing to reflect the challenges encountered in production-grade database assistant. In this work, we introduce Bird-Interact, a benchmark that restores this missing realism through: (1) a comprehensive interaction environment that couples each database with a hierarchical knowledge base, metadata files, and a function-driven user simulator, enabling models to solicit clarifications, retrieve knowledge, and recover from execution errors without human supervision; (2) two evaluation settings reflecting real-world interaction settings which contain a pre-defined conversational protocol ($c$-Interact) and a more open-ended agentic setting ($a$-Interact) in which the model autonomously decides when to query the user simulator or explore the DB environment; (3) a challenging task suite that covers the full CRUD spectrum for both business-intelligence and operational use cases, guarded by executable test cases. Each task features ambiguous and follow-up sub-tasks, requiring LLMs to engage in dynamic interaction. The suite is organized into two sets: a full set (Bird-interact-Full) of 600 tasks which unfold up to 11,796 dynamic interactions for a comprehensive overview of performance and a lite set (Bird-interact-Lite) of 300 tasks, with simplified databases for detailed behavioral analysis of interactions, and fast development of methods. Our empirical results highlight the difficulty of Bird-Interact: the most recent flagship model GPT-5 completes only 8.67% of tasks in the $c$-Interact setting and 17.00% in the $a$-Interact setting on the full task suite. Further analysis via memory grafting and Interaction Test-time Scaling (ITS) validates the importance of effective interaction for achieving success in complex, dynamic text-to-SQL tasks.
@@ 1 Introduction
Data-driven decision-making has become indispensable across modern enterprises, prompting a surge of interest in Natural Language Interfaces to Databases (NLIDB) that empower non-technical users to extract insights from relational databases using natural language (Shi et al., 2024). Motivated by this vision, a wave of methods (Pourreza et al., 2025a; b; Pourreza & Rafiei, 2023; Liu et al., 2025; Qu et al., 2024; Li et al., 2025b; Maamari et al., 2024; Sheng & Xu, 2025; Li et al., 2025a; Talaei et al., 2024; Caferoğlu & Ulusoy, 2024; Cao et al., 2024; Lee et al., 2025) based on large language models (LLMs) has recently achieved impressive text-to-SQL performance on popular single-turn benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2023b).
However, real-world data interaction is rarely a single, perfectly-formed query (Li et al., 2025c; Dinan et al., 2019). It is an iterative, stateful dialogue characterized by ambiguity (Chen et al., 2025b) and evolving goals (Wu et al., 2025). The task in Figure 1 exemplifies this complexity. To succeed, the text-to-SQL system must first engage the user to resolve the ambiguity of the term urgent care. Only with this clarified context can it generate the correct SQL. If its initial code fails an execution test, LLM must debug and revise its SQL solution based on the error feedback. After the user confirms the SQL is correct, they may proceed with a follow-up question that depends on its intermediate results. Therefore, evaluation on true practical utility LLMs with these multi-faceted aspects requires a benchmark containing a complete interactive problem-solving process, rather than isolated, single-turn SQL generation, but the entire interactive problem-solving loop.
@@
Although existing interactive text-to-SQL datasets (Yu et al., 2019b; a; Chen et al., 2025b; Guo et al., 2021; Dahl et al., 1994) have been developed, they inadequately model this reality for two primary reasons. First, most multi-turn text-to-SQL benchmarks rely on static conversation transcripts (Yu et al., 2019a; Chen et al., 2025b; Yu et al., 2019b; Guo et al., 2021). They present models with a clean interaction history without recording the failed attempts, digressions, and clarifications that occur in practice. This design may introduce a fundamental limitation: every LLM is evaluated against the same predetermined dialogue trajectory, regardless of how it would have naturally guided the interaction. This setup fails to reward intelligent interaction strategies and cannot effectively penalize conversational mess up. Second, existing benchmarks suffer from a narrow task scope, overwhelmingly focusing on read-only (SELECT-only) queries typical of business intelligence (BI) reporting. This ignores a vast and critical range of database management (DM) operations, including data manipulation (INSERT, UPDATE, DELETE), schema modifications (ALTER TABLE), and transactional control, which are also common operations in the normal DBA cycle (Chen et al., 2024).
@@ 2 Problem Definition
Each sub-task $q_{i}$ is annotated with ground-truth SQL $\sigma_{i}^{*}$ and executable test cases $\mathcal{T}_{i}$ that define correctness. A predicted solution $\sigma_{i}$ is correct if it passes all associated test cases, ensuring functional equivalence with $\sigma_{i}^{*}$. In our implementation, each task consists of two related sub-tasks ($n=2$): an initial priority sub-task $q_{1}$ containing ambiguities requiring resolution, and (2) a subsequent follow-up sub-task $q_{2}$. We evaluate system performance using: (1) Success Rate (SR): The proportion of sub-tasks completed successfully, with each sub-task scored 0 or 1. We report SR separately for sub-task 1 and sub-task 2 as an online evaluation during interaction. (2) Normalized Reward: Defined as normalized scoring according to priority weighting as designed in Appendix E to $[0,1]$ for analyzing system behaviors after interaction (offline evaluation) (Yao et al., 2022).
@@ 3.2 Interactive Task Annotation
To maintain the integrity and quality of our benchmark, we recruit 12 expert annotators through a rigorous multi-stage selection process detailed in Appendix B. We describe systematically the conversion from single-turn tasks of LiveSQLBench into multi-turn interactive scenarios through two key annotation strategies: ambiguity injection and follow-up sub-task generation:
@@
Each injected ambiguity is paired with a corresponding SQL snippet from the ground-truth query as a clarification source, which guides our user simulator in generating consistent and contextually appropriate clarifications. Quality control ensures that ambiguous queries are unsolvable without clarification yet fully reconstructable once clarifications are provided. Complete details are given in Appendix G.
@@
These follow-up sub-tasks are designed carefully using a principled 5-category taxonomy detailed in Appendix G.5. A key contribution of our benchmark is the introduction of state dependency between sub-tasks, different with other datasets (Yu et al., 2019a; b; Lee et al., 2021; Zhong et al., 2017; Li et al., 2025d). System models must reason over modified database states or the newly created objects (e.g. tables) from preceding queries to write SQLs for follow-up sub-tasks.
@@ 3.3 Function-Driven User Simulator
To ensure a more robust evaluation, we introduce a two-stage function-driven user simulator, as illustrated in Figure 3(c). In the first stage, an LLM functions as a semantic parser. It maps the system’s clarification request into one of three predefined allowed actions: AMB(), LOC(), or UNA(). AMB() is invoked for queries related to ambiguities that have been pre-annotated with the key SQL snippet. LOC() handles reasonable clarification requests that fall outside our pre-annotated ambiguities, such as questions about SQL formatting or specific sub-components. In these cases, the simulator uses an AST-based retrieval step to locate the relevant SQL fragment (detailed in Appendix M). Finally, UNA() rejects any inappropriate requests, such as attempts to elicit ground-truth answers. In the second stage, the user simulator generates a final response based on the chosen action and the annotated GT SQL with clarification source. This two-stage approach, ensures the simulator’s behavior remains predictable and controllable, while still permitting diverse and context-aware interactions. Detailed prompts are provided in Appendix Q.
@@ 3.4 Data Statistics
Table 1 reports key properties of Bird-Interact. The resulting benchmark comprises a total of 900 interactive text-to-SQL tasks, each featuring an ambiguous initial priority sub-task, dynamic clarification requirements, follow-up sub-tasks, and environmental uncertainties, collectively spanning the full CRUD spectrum (Create, Read, Update, and Delete). In Appendix D, we also conduct a comprehensive comparison against other relevant benchmarks, showing that Bird-Interact is among the most open, challenging, and long-horizon interactive benchmarks in text-to-SQL scenarios.
@@ 4 Evaluation Settings
The application of LLMs is limited by computational resources and user patience (Wen et al., 2025; Li et al., 2025e). We introduce a budget-constrained awareness mechanism to both evaluation settings, where interactions are capped by an adaptive budget and systems are informed of the remaining budget. This enables evaluation under varying budgets, including stress-testing (Ahmad et al., 2025; Hubinger, 2024) in low-budget conditions to assess the system’s ability to ask the right questions and plan effectively. The specific budget settings are detailed in the following sections.
@@ 4.1 $c$-Interact Evaluation
The budget is implemented as a constraint on the number of clarification turns. The total allowed turns, $\tau_{\text{clar}}$, are calculated as follows: $\tau_{\text{clar}}=m_{\text{amb}}+\lambda_{\text{pat}},$
@@ 4.2 $a$-Interact Evaluation
The $a$-Interact provides LLMs with autonomous planning and execution within a pre-defined action space, following ReAct paradigm (Yao et al., 2023). We model the complete database environment as a set of callable tools, containing the target database, metadata, HKB, and User Simulator, allowing the agent to determine optimal invocation strategies dynamically. In this work, we summarize and define 9 discrete actions common to text-to-SQL with details in Appendix I. Bird-Interact also supports customized scaffolds, details can be found in Appendix I.2.
@@
To reflect the varying computational costs of different actions, we implement a budget-constrained evaluation framework where each action consumes a predetermined amount of budget, encouraging cost-effective action sequences. The total budget for each task is $B=B_{\text{base}}+2\,m_{\text{amb}}+2\,\lambda_{\text{pat}}$, where $B_{\text{base}}=6$ is the base budget, $m_{\text{amb}}$ is the number of annotated ambiguity points, and $\lambda_{\text{pat}}$ is the user patience parameter, maintaining consistency with the $c$-Interact framework. This setting evaluates the agent’s ability to achieve high performance under resource constraints while balancing thoroughness with efficiency. Further details of action costs are provided in Appendix I.
@@ 5 Experiment
We benchmark 7 recent and powerful LLMs (2 open-source, 5 closed-source) as system models via a fresh PostgreSQL 14 Docker instance for more stable evaluation. We set the user patience to 3 by default and $a$-Interact base budget of 6. All models use temperature=0 and top_p=1, with default reasoning settings, conducting single runs due to cost (full details in Appendix H.2 and H.3).
@@
Model Priority Question (Success Rate %) $\uparrow$ Follow Ups (Success Rate %) $\uparrow$ Reward* $\uparrow$
Avg.
Cost $\downarrow$
@@ 5.1 Main Results
Bird-Interact remains challenging, leaving ample room for future improvement. Even the strongest models in our study, GPT-5 and Gemini-2.5-Pro, capture only 20.92% and 25.52% of the available reward respectively, in the $c$-Interact and $a$-Interact mode. Absolute success rates reveal similar limitations: no more than 16.33% of tasks are solved end-to-end in $c$-Interact and 17.00% in $a$-Interact, with most models falling in substantially lower rates.
@@ 5.2 Interaction Analysis
A notable finding is the underperformance of the flagship model, GPT-5, on the $c$-Interact, despite its strong performance on many single-turn tasks (Phan et al., 2025; Glazer et al., 2024; Rein et al., 2024). Therefore, we hypothesize that this stems from a deficiency in its interactive communication abilities rather than its core generation capability. To test this hypothesis, we conduct an experiment termed Memory Grafting. In this setup, we provide GPT-5 with the ambiguity resolution histories from two other better models, Qwen-3-Coder and O3-mini, before asking it to generate the final SQL query. The results, presented in Figure 5, show that GPT-5’s performance improves significantly when leveraging the interaction history from either model. This finding indicates that while GPT-5 possesses robust SQL generation capabilities, a more effective communication schema is required to help it achieve satisfactory outcomes for user tasks. We also further analyze the patterns for effective communication in Appendix O.
@@
We analyze action distributions across 7 system models and find concentration in two primary actions: submit (direct code execution with error feedback) and ask (user clarification requests), which together comprise 60.87% of all actions. Despite being the most computationally expensive actions (Figure 3), models favor these over systematic exploration behaviors like knowledge and schema retrieval. This suggests LLMs prefer direct trial-and-error execution over comprehensive environment exploration, likely due to pre-training biases. Future work should incentivize broader tool utilization for complex interactive tasks. Additional analysis on the Full set appears in Appendix I.
@@ 6 User Simulator Analysis
To provide an objective and comprehensive observation of different user simulator mechanisms, we construct a static dataset called UserSim-Guard, comprising 1,989 questions with reference actions labeled by human experts. Detailed information regarding the distribution and annotation procedures can be found in Appendix N. We employed an LLM-as-Judge (Zheng et al., 2023) evaluation framework using Qwen-2.5-72B and Llama-3.1-70B as independent evaluators to mitigate potential self-enhancement bias. Our analysis reveals significant reliability concerns with conventional user simulator designs. Specifically, as shown in Figure 6, when confronted with Unanswerable (UNA) questions, baseline user simulators consistently fail to implement safeguards, resulting in unfair or inappropriate feedback generation with over 34% failure rate. In contrast, our proposed function-driven approach demonstrates substantially improved reliability, with only 5.9% of responses falling into problematic categories. This represents a significant improvement in user simulator robustness and reliability compared to baseline approaches.
@@
We evaluate alignment between our user simulators and actual human behavior by having human experts interact with 7 system models on 100 randomly sampled tasks across BI and DM domains. We then compute correlations (Ivey et al., 2024; Kong et al., 2024) between success rates (SR) achieved by human users versus our simulators across the same tasks. As shown in Table 3, function-driven simulators demonstrate significantly stronger alignment with human behavior: GPT-4o with function calling achieves 0.84 Pearson correlation (p = 0.02) compared to 0.61 without function calling (p = 0.14), while Gemini-2.0-Flash shows similar improvements. These results confirm that incorporating our designed mechanism produces more realistic user simulators that better reflect actual human-AI interaction patterns (detailed analysis in Appendix N).

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance on single-turn text-to-SQL tasks, but real-world database applications predominantly require multi-turn interactions to handle ambiguous queries, execution errors, and evolving user requirements. Existing multi-turn benchmarks fall short of capturing this complexity, either by treating conversation histories as static context or by limiting evaluation to narrow, read-only (SELECT-ONLY) operations, thereby potentially failing to reflect the challenges encountered in production-grade database assistant. In this work, we introduce Bird-Interact, a benchmark that restores this missing realism through: (1) a comprehensive interaction environment that couples each database with a hierarchical knowledge base, metadata files, and a function-driven user simulator, enabling models to solicit clarifications, retrieve knowledge, and recover from execution errors without human supervision; (2) two evaluation settings reflecting real-world interaction settings which contain a pre-defined conversational protocol ($c$-Interact) and a more open-ended agentic setting ($a$-Interact) in which the model autonomously decides when to query the user simulator or explore the DB environment; (3) a challenging task suite that covers the full CRUD spectrum for both business-intelligence and operational use cases, guarded by executable test cases. Each task features ambiguous and follow-up sub-tasks, requiring LLMs to engage in dynamic interaction. The suite is organized into two sets: a full set (Bird-interact-Full) of 600 tasks which unfold up to 11,796 dynamic interactions for a comprehensive overview of performance and a lite set (Bird-interact-Lite) of 300 tasks, with simplified databases for detailed behavioral analysis of interactions, and fast development of methods. Our empirical results highlight the difficulty of Bird-Interact: the most recent flagship model GPT-5 completes only 8.67% of tasks in the $c$-Interact setting and 17.00% in the $a$-Interact setting on the full task suite. Further analysis via memory grafting and Interaction Test-time Scaling (ITS) validates the importance of effective interaction for achieving success in dynamic text-to-SQL tasks.
@@ 1 Introduction
Data-driven decision-making has become indispensable across modern enterprises, prompting a surge of interest in Natural Language Interfaces to Databases (NLIDB) that empower non-technical users to extract insights from relational databases using natural language (Shi et al., 2024). Motivated by this vision, a wave of methods (Pourreza et al., 2025a; b; Pourreza & Rafiei, 2023; Liu et al., 2025; Qu et al., 2024; Li et al., 2025b; Maamari et al., 2024; Sheng & Shuai, 2025; Li et al., 2025a; Talaei et al., 2024; Caferoğlu & Ulusoy, 2024; Cao et al., 2024; Lee et al., 2025) based on large language models (LLMs) has recently achieved impressive text-to-SQL performance on popular single-turn benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2023b).
However, real-world data interaction is rarely a single, perfectly-formed query (Li et al., 2025c; Dinan et al., 2019). It is an iterative, stateful dialogue characterized by ambiguity (Chen et al., 2025b) and evolving goals (Wu et al., 2025). The task in Figure 1 exemplifies this complexity. To succeed, the text-to-SQL system must first engage the user to resolve the ambiguity of the term urgent care. Only with this clarified context can it generate the correct SQL. If its initial code fails an execution test, LLM must debug and revise its SQL solution based on the error feedback. After the user confirms the SQL is correct, they may proceed with a follow-up question that depends on its intermediate results. Therefore, evaluation on true practical utility LLMs with these multi-faceted aspects requires a benchmark containing a complete interactive problem-solving process, rather than isolated, single-turn SQL generation.
@@
Although existing interactive text-to-SQL datasets (Yu et al., 2019b; a; Chen et al., 2025b; Guo et al., 2021; Dahl et al., ) have been developed, they inadequately model this reality for two primary reasons. First, most multi-turn text-to-SQL benchmarks rely on static conversation transcripts (Yu et al., 2019a; Chen et al., 2025b; Yu et al., 2019b; Guo et al., 2021). They present models with a clean interaction history without recording the failed attempts, digressions, and clarifications that occur in practice. This design may introduce a fundamental limitation: every LLM is evaluated against the same predetermined dialogue trajectory, regardless of how it would have naturally guided the interaction. This setup fails to reward intelligent interaction strategies and cannot effectively penalize conversational mess up. Second, existing benchmarks suffer from a narrow task scope, focusing on read-only (SELECT-only) queries typical of business intelligence (BI) reporting. This ignores a vast and critical range of database management (DM) operations, including data manipulation (INSERT, UPDATE, DELETE), schema modifications (ALTER TABLE), and transactional control, which are also common operations in the normal DBA cycle (Chen et al., 2024).
@@ 2 Problem Definition
Each sub-task $q_{i}$ is annotated with ground-truth SQL $\sigma_{i}^{*}$ and executable test cases $\mathcal{T}_{i}$ that define correctness. A predicted solution $\sigma_{i}$ is correct if it passes all associated test cases, ensuring functional equivalence with $\sigma_{i}^{*}$. In our implementation, each task consists of two related sub-tasks ($n=2$): an initial priority sub-task $q_{1}$ containing ambiguities requiring resolution, and (2) a subsequent follow-up sub-task $q_{2}$. We evaluate system performance using: (1) Success Rate (SR): The proportion of sub-tasks completed successfully, with each sub-task scored 0 or 1. We report SR separately for sub-task 1 and sub-task 2 as an online evaluation during interaction. (2) Normalized Reward: Defined as normalized scoring according to priority weighting as designed in Appendix F to $[0,1]$ for analyzing system behaviors after interaction (offline evaluation) (Yao et al., 2022).
@@ 3.2 Interactive Task Annotation
To maintain the integrity and quality of our benchmark, we recruit 12 expert annotators through a rigorous multi-stage selection process detailed in Appendix C. We describe systematically the conversion from single-turn tasks of LiveSQLBench into multi-turn interactive scenarios through two key annotation strategies: ambiguity injection and follow-up sub-task generation:
@@
Each injected ambiguity is paired with a corresponding SQL snippet from the ground-truth query as a clarification source, which guides our user simulator in generating consistent and contextually appropriate clarifications. Quality control ensures that ambiguous queries are unsolvable without clarification yet fully reconstructable once clarifications are provided. Complete details are given in Appendix H.
@@
These follow-up sub-tasks are designed carefully using a principled 5-category taxonomy detailed in Appendix H.5. A key contribution of our benchmark is the introduction of state dependency between sub-tasks, different with other datasets (Yu et al., 2019a; b; Lee et al., 2021; Zhong et al., 2017; Li et al., 2025d). System models must reason over modified database states or the newly created objects (e.g. tables) from preceding queries to write SQLs for follow-up sub-tasks.
@@ 3.3 Function-Driven User Simulator
To ensure a more robust evaluation, we introduce a two-stage function-driven user simulator, as illustrated in Figure 3(c). In the first stage, an LLM functions as a semantic parser. It maps the system’s clarification request into one of three predefined allowed actions: AMB(), LOC(), or UNA(). AMB() is invoked for queries related to ambiguities that have been pre-annotated with the key SQL snippet. LOC() handles reasonable clarification requests that fall outside our pre-annotated ambiguities, such as questions about SQL formatting or specific sub-components. In these cases, the simulator uses an AST-based retrieval step to locate the relevant SQL fragment (detailed in Appendix N). Finally, UNA() rejects any inappropriate requests, such as attempts to elicit ground-truth answers. In the second stage, the user simulator generates a final response based on the chosen action and the annotated GT SQL with clarification source. This two-stage approach ensures the simulator’s behavior remains predictable and controllable, while still permitting diverse and context-aware interactions. Detailed prompts are provided in Appendix R.
@@ 3.4 Data Statistics
Table 1 reports key properties of Bird-Interact. The resulting benchmark comprises a total of 900 interactive text-to-SQL tasks, each featuring an ambiguous initial priority sub-task, dynamic clarification requirements, follow-up sub-tasks, and environmental uncertainties, collectively spanning the full CRUD spectrum (Create, Read, Update, and Delete). In Appendix E, we also conduct a comprehensive comparison against other relevant benchmarks, showing that Bird-Interact is among the most open, challenging, and long-horizon interactive benchmarks in text-to-SQL scenarios.
@@ 4 Evaluation Settings
The application of LLMs is limited by computational resources and user patience (Wen et al., 2025; Li et al., 2025e). We introduce a budget-constrained awareness mechanism to both evaluation settings, where interactions are capped by an adaptive budget and systems are informed of the remaining budget. This enables evaluation under varying budgets, including stress-testing (Ahmad et al., 2025; Zhang et al., 2025) in low-budget conditions to assess the system’s ability to ask the right questions and plan effectively. The specific budget settings are detailed in the following sections.
@@ 4.1 $c$-Interact Evaluation
The budget is implemented as a constraint on the number of clarification turns. The total allowed turns, $\tau_{\text{clar}}$, are calculated as follows: $\tau_{\text{clar}}=m_{\text{amb}}+\lambda_{\text{pat}}.$
@@ 4.2 $a$-Interact Evaluation
The $a$-Interact provides LLMs with autonomous planning and execution within a pre-defined action space, following ReAct paradigm (Yao et al., 2023). We model the complete database environment as a set of callable tools, containing the target database, metadata, HKB, and User Simulator, allowing the agent to determine optimal invocation strategies dynamically. In this work, we summarize and define 9 discrete actions common to text-to-SQL with details in Appendix J. Bird-Interact also supports customized scaffolds, details can be found in Appendix J.2.
@@
To reflect the varying computational costs of different actions, we implement a budget-constrained evaluation framework where each action consumes a predetermined amount of budget, encouraging cost-effective action sequences. The total budget for each task is $B=B_{\text{base}}+2\,m_{\text{amb}}+2\,\lambda_{\text{pat}}$, where $B_{\text{base}}=6$ is the base budget, $m_{\text{amb}}$ is the number of annotated ambiguity points, and $\lambda_{\text{pat}}$ is the user patience parameter, maintaining consistency with the $c$-Interact framework. This setting evaluates the agent’s ability to achieve high performance under resource constraints while balancing thoroughness with efficiency. Further details of action costs are provided in Appendix J.
@@ 5 Experiment
We benchmark 7 recent and powerful LLMs (2 open-source, 5 closed-source) as system models via a fresh PostgreSQL 14 Docker instance for more stable evaluation. We set the user patience to 3 by default and $a$-Interact base budget of 6. All models use temperature=0 and top_p=1, with default reasoning settings, conducting single runs due to cost (full details in Appendix I.2 and I.3).
@@
Model Priority Question (Success Rate %) $\uparrow$ Follow Ups (Success Rate %) $\uparrow$ Reward* $\uparrow$ Avg. Cost $\downarrow$
@@ 5.1 Main Results
Bird-Interact Remains Challenging, Leaving Ample Room for Future Improvement. Even the strongest models in our study, Gemini-2.5-Pro and GPT-5, capture only 20.92% and 25.52% of the available reward respectively, in the $c$-Interact and $a$-Interact mode. Absolute success rates reveal similar limitations: no more than 16.33% of tasks are solved end-to-end in $c$-Interact and 17.00% in $a$-Interact, with most models falling in substantially lower rates.
@@ 5.2 Interaction Analysis
A notable finding is the underperformance of the flagship model, GPT-5, on the $c$-Interact, despite its strong performance on many single-turn tasks (Phan et al., 2025; Glazer et al., 2024; Rein et al., 2024). Therefore, we hypothesize that this stems from a deficiency in its interactive communication abilities rather than its core generation capability. To test this hypothesis, we conduct an experiment termed Memory Grafting. In this setup, we provide GPT-5 with the ambiguity resolution histories from two other better models, Qwen-3-Coder and O3-mini, before asking it to generate the final SQL query. The results, presented in Figure 5, show that GPT-5’s performance improves significantly when leveraging the interaction history from either model. This finding indicates that while GPT-5 possesses robust SQL generation capabilities, a more effective communication schema is required to help it achieve satisfactory outcomes for user tasks. We also further analyze the patterns for effective communication in Appendix P.
@@
We analyze action distributions across 7 system models and find concentration in two primary actions: submit (direct code execution with error feedback) and ask (user clarification requests), which together comprise 60.87% of all actions. Despite being the most computationally expensive actions (Figure 3), models favor these over systematic exploration behaviors like knowledge and schema retrieval. This suggests LLMs prefer direct trial-and-error execution over comprehensive environment exploration, likely due to pre-training biases. Future work should incentivize broader tool utilization for complex interactive tasks. Additional analysis on the Full set appears in Appendix J.
@@ 6 User Simulator Analysis
To provide an objective and comprehensive observation of different user simulator mechanisms, we construct a static dataset called UserSim-Guard, comprising 2,100 questions with reference actions labeled by human experts. Detailed information regarding the distribution and annotation procedures can be found in Appendix O. We employed an LLM-as-Judge (Zheng et al., 2023) evaluation framework using Qwen3-235B-A22B-Instruct-2507 as independent evaluators. Our analysis reveals significant reliability concerns with conventional user simulator designs. Specifically, as shown in Figure 6, when confronted with Unanswerable (UNA) questions, baseline user simulators consistently fail to implement safeguards, resulting in unfair or inappropriate feedback generation with failure rates reaching up to 67.4% depending on the backbone model. In contrast, our proposed function-driven approach demonstrates substantially improved reliability, reducing the failure rate to as low as 2.7%. This represents a dramatic improvement in user simulator robustness and reliability compared to baseline approaches.
@@
We evaluate alignment between our user simulators and actual human behavior by having human experts interact with 7 system models on 100 randomly sampled tasks across BI and DM domains. We then compute correlations (Ivey et al., 2024; Kong et al., 2024) between success rates (SR) achieved by human users versus our simulators across the same tasks. As shown in Table 3, function-driven simulators demonstrate significantly stronger alignment with human behavior: GPT-4o with function calling achieves 0.84 Pearson correlation (p = 0.02) compared to 0.61 without function calling (p = 0.14), while Gemini-2.0-Flash shows similar improvements. These results confirm that incorporating our designed mechanism produces more realistic user simulators that better reflect actual human-AI interaction patterns (detailed analysis in Appendix O).